OpenStudy (anonymous):

Perform the indicated operation. Simplify if possible. \[ \frac{ 2x }{ x^2-5x+6 }- \frac{ 6 }{ x^2-5x+6 }\]

terenzreignz (terenzreignz):

Well, they have the same denominator, so you can just go ahead and subtract their numerators (and place the same denominator under them) \[\Large \frac{a}c-\frac{b}c = \frac{a-b}c\]

OpenStudy (anonymous):

\[\frac{ 2(x+3 }{ (x+3)(x_3) }\] so it would be this

terenzreignz (terenzreignz):

Step by step, it should be \[\Large \frac{2x - 6}{x^2 - 5x +6}\]

terenzreignz (terenzreignz):

Work from here, factor out what can be factored, and cancel what may be cancelled... In particular, factor the denominator...